# Theory of Mind for Explainable Human-Robot Interaction

### v2

@@ Introduction
As interactions between humans and robots become increasingly common (Lee-service-robot), it is intuitive to seek more human-like modes of interaction to be able to understand robots’ behaviors (sridharan-2019; Kerzel2023). This need naturally motivates the application of ToM in HRI. ToM refers to the human ability to attribute mental states such as beliefs, desires, and intentions to oneself and others to predict and explain behavior (Premack_Woodruff_1978). When embedded in robots, ToM methods emphasize understanding and adapting to users’ mental states, and this can be used to produce explanations that are often more intuitive and user-friendly (williams-2022). ToM also allows robots to interpret and respond to users’ inferred mental states, fostering more natural, adaptive, and transparent interactions (Yuan). On the other hand, XAI aims to make black box models more transparent and interpretable; however, it frequently overlooks user-centered evaluations (Rong).
@@
Several studies examine whether humans naturally attribute ToM to robots even in the absence of explicit ToM mechanisms. A first study found that humans are able to interpret robots’ behavior similarly to human behavior, provided that the robots display distinct and interpretable social cues. However, when a robot’s cues deviate from human expectations, this understanding diminishes (Bank). A second study, which examined the robustness and conviction of large language models (LLMs), demonstrated that while LLMs can serve as a useful tool in human–robot interaction (Becker), they do not function as reliable ToM agents (Verma). These findings suggest that effective human–robot interaction is facilitated when robots produce responses that align with typical human behavior, and highlight the need for the integration of explicit ToM-like mechanisms in robotic systems.
@@
Bank $\checkmark$ $\checkmark$ $\checkmark$
Mou $\checkmark$ $\checkmark$
Cantucci $\checkmark$ $\checkmark$ $\checkmark$
Kerzel $\checkmark$ $\checkmark$ $\checkmark$
Shvo $\checkmark$ $\checkmark$ $\checkmark$
Yuan $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Verma $\checkmark$ $\checkmark$ $\checkmark$
Angelopoulos $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
A second line of research has investigated embedding ToM-like reasoning directly within robots and assessing its impact on trust, helpfulness, and mutual understanding. Some studies have focused on evaluating user perception, revealing that robots equipped with ToM capabilities are perceived more positively (Mou), particularly when they provide assistance aligned with users’ goals (Cantucci). Similarly, robots that reason about human beliefs are generally considered more helpful and socially competent (Shvo), and are also regarded as more trustworthy (Angelopoulos). At the same time, when providing explanations, robots may fail to enhance user understanding or improve decision-making, as not all explanations are equally effective (Yuan). In contrast, approaches that implement multiple levels of explanation have been shown to improve user comprehension and the interaction (Kerzel). Although these studies evaluate human–AI collaboration and occasionally describe their work as XAI, none assess it using XAI-specific criteria. Moreover, none have explicitly integrated ToM with XAI, highlighting a gap that our work addresses.
@@ A Comprehensive Evaluation of ToM
While the field of ToM primarily claims to enhance user understanding, trust, and, more broadly, human–AI collaboration, these claims are often not systematically evaluated. This lack of evaluation stems from the fact that, if ToM purports to provide explanations for users, it should be assessed using the same criteria applied in XAI. Indeed, these objectives align closely with those of XAI, which aims to design AI systems that are interpretable and comprehensible to humans (Rong). We therefore propose to systematically evaluate ToM in line with the rigor of its claims using an XAI evaluation framework.
From this perspective, we examined whether existing ToM studies could be assessed using an XAI framework, positioning ToM itself as a framework for explanations. To this end, we evaluated state-of-the-art ToM studies using the seven desiderata defined in the eValuation XAI (VXAI) framework (dembinsky), which integrates the principal evaluation criteria identified in recent reviews and provides a comprehensive and systematic approach to XAI evaluation. We evaluated the studies based on whether they explicitly addressed each of the VXAI desiderata in their work. The criteria and detailed definitions of the desiderata are provided in AppendixA.
@@
Effective XAI should provide explanations that are both faithful to the system’s reasoning and understandable to users (Rong). Since existing methods often overlook user-centered design (Rong), we propose an integrated approach that combines ToM’s user focus with model-centered XAI techniques. This integration enables explanations that are simultaneously interpretable for users and faithful to the model’s internal reasoning. While some existing ToM integrations have proposed using Bayesian reinforcement learning to model user behavior, future work could explore the use of behavior trees or explainable reinforcement learning (XRL) within ToM-based systems. Such approaches may support adaptive reasoning while further enhancing the fidelity of explanations.
@@ Appendix A Appendix 1
The seven desiderata proposed in VXAI (dembinsky) can be described as follows:

### v3

@@ Introduction
As interactions between humans and robots become increasingly common (Lee 2021), it is intuitive to seek more human-like modes of interaction to be able to understand robots’ behaviors (Sridharan and Meadows 2019; Kerzel et al. 2023). This need naturally motivates the application of ToM in HRI. ToM refers to the human ability to attribute mental states such as beliefs, desires, and intentions to oneself and others to predict and explain behavior (Premack and Woodruff 1978). When embedded in robots, ToM methods emphasize understanding and adapting to users’ mental states, and this can be used to produce explanations that are often more intuitive and user-friendly (Williams et al. 2022). ToM also allows robots to interpret and respond to users’ inferred mental states, fostering more natural, adaptive, and transparent interactions (Yuan et al. 2022). On the other hand, XAI aims to make black box models more transparent and interpretable; however, it frequently overlooks user-centered evaluations (Rong et al. 2024).
@@
Several studies examine whether humans naturally attribute ToM to robots even in the absence of explicit ToM mechanisms. A first study found that humans are able to interpret robots’ behavior similarly to human behavior, provided that the robots display distinct and interpretable social cues. However, when a robot’s cues deviate from human expectations, this understanding diminishes (Banks,Jaime 2020). A second study, which examined the robustness and conviction of large language models (LLMs), demonstrated that while LLMs can serve as a useful tool in human–robot interaction (Becker et al. 2025), they do not function as reliable ToM agents (Verma et al. 2024). These findings suggest that effective human–robot interaction is facilitated when robots produce responses that align with typical human behavior, and highlight the need for the integration of explicit ToM-like mechanisms in robotic systems.
@@
Banks,Jaime (2020) $\checkmark$ $\checkmark$ $\checkmark$
Mou et al. (2020) $\checkmark$ $\checkmark$
Cantucci and Falcone (2022) $\checkmark$ $\checkmark$ $\checkmark$
Kerzel et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$
Shvo et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$
Yuan et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Verma et al. (2024) $\checkmark$ $\checkmark$ $\checkmark$
Angelopoulos et al. (2025) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
A second line of research has investigated embedding ToM-like reasoning directly within robots and assessing its impact on trust, helpfulness, and mutual understanding. Some studies have focused on evaluating user perception, revealing that robots equipped with ToM capabilities are perceived more positively (Mou et al. 2020), particularly when they provide assistance aligned with users’ goals (Cantucci and Falcone 2022). Similarly, robots that reason about human beliefs are generally considered more helpful and socially competent (Shvo et al. 2022), and are also regarded as more trustworthy (Angelopoulos et al. 2025). At the same time, when providing explanations, robots may fail to enhance user understanding or improve decision-making, as not all explanations are equally effective (Yuan et al. 2022). In contrast, approaches that implement multiple levels of explanation have been shown to improve user comprehension and the interaction (Kerzel et al. 2022). Although these studies evaluate human–AI collaboration and occasionally describe their work as XAI, none assess it using XAI-specific criteria. Moreover, none have explicitly integrated ToM with XAI, highlighting a gap that our work addresses.
@@ A Comprehensive Evaluation of ToM
While the field of ToM primarily claims to enhance user understanding, trust, and, more broadly, human–AI collaboration, these claims are often not systematically evaluated. This lack of evaluation stems from the fact that, if ToM purports to provide explanations for users, it should be assessed using the same criteria applied in XAI. Indeed, these objectives align closely with those of XAI, which aims to design AI systems that are interpretable and comprehensible to humans (Rong et al. 2024). We therefore propose to systematically evaluate ToM in line with the rigor of its claims using an XAI evaluation framework.
From this perspective, we examined whether existing ToM studies could be assessed using an XAI framework, positioning ToM itself as a framework for explanations. To this end, we evaluated state-of-the-art ToM studies using the seven desiderata defined in the eValuation XAI (VXAI) framework (Dembinsky et al. 2025), which integrates the principal evaluation criteria identified in recent reviews and provides a comprehensive and systematic approach to XAI evaluation. We evaluated the studies based on whether they explicitly addressed each of the VXAI desiderata in their work. The criteria and detailed definitions of the desiderata are provided in AppendixA.
@@
Effective XAI should provide explanations that are both faithful to the system’s reasoning and understandable to users (Rong et al. 2024). Since existing methods often overlook user-centered design (Rong et al. 2024), we propose an integrated approach that combines ToM’s user focus with model-centered XAI techniques. This integration enables explanations that are simultaneously interpretable for users and faithful to the model’s internal reasoning. While some existing ToM integrations have proposed using Bayesian reinforcement learning to model user behavior, future work could explore the use of behavior trees or explainable reinforcement learning (XRL) within ToM-based systems. Such approaches may support adaptive reasoning while further enhancing the fidelity of explanations.
@@ Appendix A Appendix 1
The seven desiderata proposed in VXAI (Dembinsky et al. 2025) can be described as follows: